OpenStudy (mendicant_bias):

I'm having trouble figuring out how to evaluate a particularly ugly inverse laplace form and need help; posted below in a moment.

OpenStudy (mendicant_bias):

\[L^{-1}\left\{Y(s)\right\}= L^{-1} \left\{\vphantom{} \frac{s+3}{[(s+3)^{2}+4](s+1)} \right\}\] Any immediate ideas? Because I can see this getting ugly real quick. My first thought is to try to expand the denominator and then recombine it as a different set of factors possibly?

OpenStudy (sidsiddhartha):

are u allowed to use convolution ?

OpenStudy (sidsiddhartha):

it will be very easy if we use it take two separe terms\[F_1(s)=\frac{ s+3 }{ (s+3)^2 +4}\\and\\ F_2(s)=\frac{ 1 }{ s+1 }\\\]

OpenStudy (mendicant_bias):

Not allowed to use convolution, hee

OpenStudy (mendicant_bias):

Very last problem in the section one prior XP

OpenStudy (sidsiddhartha):

now take the inverse laplace of the first function\[f_1(t)=L^{-1}(F_1(s))=L^{-1}\frac{ s+3 }{ (s+3)^2+4 }=e^{-3t}*L^{-1}\frac{ s }{ s^2+4 }\\=e^{-3t}.\cos(2t)\]

OpenStudy (sidsiddhartha):

\[f_2(t)=L^{-1}\frac{ 1 }{ s+1 }=sint\\now\\finally ~~f(t)=\int\limits_{0}^{t}f_1(u).f_2(t-u)du\] this will make it a lot easier and if u're not allowed then laborious partial fraction \[\frac{ s+3 }{ (s^2+6s+13)(s+1) }=\frac{ A }{ s+1 }+\frac{ Bs+C }{ s^2+6s+13 }\] this way-----

OpenStudy (mendicant_bias):

Alright, so the partial fraction approach is to bust it up, thank you so much. Just was making sure I wasn't missing something. I also promise I'll do those convolution problems you posted!

OpenStudy (sidsiddhartha):

yes try that :) sole purpose of convolution is to reduce the complexity